Kagiso wants to buy a new gaming computer for R40 000. He decides to save by depositing an amount of R400 quarterly into an account earning 16% interest per year, compounded quarterly. The approximate number of quarters it will take Kagiso to have R40 000 available is


1
Expert's answer
2022-03-19T02:41:44-0400

Solution


Interest rate "=16\%" per year "\frac{16\%}{4}=4\%" per quarter


Future Value "FV = R 40,000"


Deposit Per Quarter "A = R 400"


The number of periods required to generate "R 40,000" is "n =?"


"\begin{gathered}
 n = \ln \left( {1 + \frac{{\left( {FV \times r} \right)}}{A}} \right) \div \ln \left( {1 + r} \right) \\
 n = \ln \left( {1 + \frac{{\left( {40,000 \times \frac{4}{{100}}} \right)}}{{400}}} \right) \div \ln \left( {1 + \frac{4}{{100}}} \right) \\
 n = 41.03540663 \\ 
\end{gathered}"


Hence it will Kagiso take around 41 quarters to get the amount of "R 40,000"


The same has been displayed by the following table